OpenStudy (anonymous):

Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately what maximum height does the soccer ball reach? A. 0.8 ft B. 1.6 ft C. 9.8 ft D. 19.6 ft I believe the answer is C? Is that correct?

OpenStudy (anonymous):

Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately, how long does it take until the soccer ball hits the ground again? A. 0.6 sec B. 0.8 sec C. 1.6 secs D. 2.8 secs I believe the answer for this one is B? Is that correct?

OpenStudy (turingtest):

this is physics so please post in the physics section in the future this time I will give you the relevent formula\[v_f^2=v_i^2+2ad\]which in the case of free fall, at the top becomes\[0=v_i^2-2gh\]solve for h to get the first answer for the second yuo will need to post in the proper section

OpenStudy (anonymous):

I have some more questions here, if you woldn't mind chekcing them out: http://answers.yahoo.com/question/index?qid=20120406175748AARkAqy

OpenStudy (turingtest):

This site is about teaching ask each of those questions individually in the proper sections and you will get more help gotta eat now. later

OpenStudy (anonymous):

This is for the first one h'(t) = -32t + 25 0 = -32t + 25 t = 25/32 Plug it back into the original equation. h(t) = -16t^2 + 25t + 0 h(25/32) = -16(25/32)^2 + 25(25/32) + 0 h(25/32) = -9.766 + 19.531 h(25/32) = 9.765

OpenStudy (turingtest):

looks good to me, but I only think in metric

OpenStudy (turingtest):

so I'll trust that g=32ft/s^2 lol

OpenStudy (anonymous):

That's how I found the answer, I was just checking it, this is actually for an at home study quiz for algebra.... lol

OpenStudy (anonymous):

Physics questions on an algebra quiz, lol

OpenStudy (turingtest):

my solution is algebraic, yours uses calculus they are both valid gotta go, good luck :D

OpenStudy (anonymous):

okay, thanks for your help! :)